OpenStudy (anonymous):

If dy/dx = y cos x and y = 3 when x =0, then y=...?

OpenStudy (anonymous):

Please explain what i need to do becuase im not even sure how to approach this problem

OpenStudy (anonymous):

That is exactly how the question is asked: If dy/dx = y cos x and y = 3 when x = 0, then y =....? Here are the answer choices a) e^(sinx) +2 b) e^(sinx) + 3 c) sin x + 3 d) sin x + 3e^x e) 3e^ (sinx)

OpenStudy (anonymous):

Does teh question make more sense now

OpenStudy (amistre64):

it looks rather seperable to me

OpenStudy (amistre64):

\[\frac{dy}{dx}=ycos(x)\] \[\int \left(\frac{1}{y}dy=cos(x)dx\right)\]

OpenStudy (amistre64):

\[ln(y)=sin(x)+c\] \[y = Ce^{sin(x)}\]

OpenStudy (anonymous):

wait how did u get taht integral

OpenStudy (amistre64):

multiply it all by dx/y

OpenStudy (amistre64):

basically, your collecting all the y terms to one side and all the x terms to the other and integrating the whole shebang

OpenStudy (anonymous):

ok i get that so now hoe did u get y=Ce^sinx

OpenStudy (amistre64):

in order to undo the ln; you have to e^ it all

OpenStudy (amistre64):

e^ln(y) = y e^(sin(x)+c) = e^sin(x) e^c e^c is just some arbitrary constnat still; e^c = C y = Ce^sin(x)

OpenStudy (anonymous):

ok so now how do we find out wat that constant is b/c 2 of the asnwers have 2 and 3 as constants

OpenStudy (amistre64):

yeah, they all have something added; look at the results we got; anything added?

OpenStudy (anonymous):

oh it has to be e) 3e^sinx becasue thats teh only one with a cosntant being multiplied

OpenStudy (amistre64):

thats the short of it yes; or we could test that by making x=0; sin(x)= 0 e^0 = 1; and 3 = C*1; C=3 y = 3e^sin(x)

OpenStudy (anonymous):

THANKS ALOT =)

OpenStudy (amistre64):

yw

OpenStudy (rea201):

Is this right

OpenStudy (amistre64):

what about the detailed explanations are you confused about?

OpenStudy (rea201):

Oh I looked this over a couple more times and understood. I got confused on the step when ln(y)=sinx+C changed to y=Ce^sinx

OpenStudy (amistre64):

it makes sense now tho? i just used exponent properties to condense it is all

OpenStudy (rea201):

Yes it makes sense. You worked it out really well